# Language on Demand, Knowledge at Core: Composing LLMs with Encoder-Decoder Translation Models for Extensible Multilinguality

### 18 Mar 2026 v1

@@ 2.1 Data-Level Multilingual Enhancement for LLMs
A line of work augments the multilingual capabilities of LLMs at the data level by constructing multilingual training corpora using pretrained multilingual or machine translation models Li et al. (2023); Zhang et al. (2023, 2024a, 2024b). Typical approaches translate English instruction or task data into multiple languages Chen et al. (2024), or employ translation-based prompting schemes that map non-English inputs into English before task execution Qin et al. (2023); Chai et al. (2025). These methods are widely adopted for multilingual instruction tuning and cross-lingual transfer. Such approaches generally require continual multilingual training of LLMs, which may introduce translation noise and interfere with existing language capabilities. In practice, balancing performance across high- and low-resource languages remains challenging, as gains on low-resource languages often come at degradation on high-resource ones Gao et al. (2024). In contrast, XBridge achieves multilingual generalization through model composition without multilingual retraining of the LLM.
@@ 3.1 Architecture
Formally, given an input sequence $\mathbf{x}=(x_{1},\ldots,x_{n})$ in language $L_{x}$, we first encode it with the pretrained multilingual encoder $\mathtt{Enc}(\cdot)$, producing contextual representations $\mathbf{H}^{\mathbf{x}}\in\mathbb{R}^{n\times d_{e}}$. To bridge the representation gap between the multilingual encoder and LLM, we apply a lightweight mapping $\mathtt{Mapping_{enc}}(\cdot)$ that projects $\mathbf{H}^{\mathbf{x}}$ into the LLM representation space, yielding $\tilde{\mathbf{H}}^{\mathbf{x}}\in\mathbb{R}^{n\times d_{l}}$. The mapped encoder representations are then injected into the LLM together with a high-resource (English) instruction prompt, enabling the LLM to perform general knowledge processing conditioned on encoder semantics. Let $\mathbf{z}=(z_{1},\ldots,z_{m})$ denote the sequence of English tokens generated by the LLM. Rather than using the final-layer hidden states, we extract the penultimate-layer hidden states, denoted as $\mathbf{H}^{\mathbf{z^{\prime}}}\in\mathbb{R}^{m\times d_{l}}$, as Zhang et al. (2025) show that the last layer is often tightly aligned with the output vocabulary space, while non-final layers retain richer semantic information.
@@ 5.2 Generalization to Untuned Languages
To examine whether the cross-model mappings learned by XBridge are language-agnostic rather than simply tied to specific training languages, we evaluate cross-model cross-lingual transfer on 42 untuned languages (listed in Table 3) in Figure 5, based on variants of Section 5.1.
@@ 5.6 Impact of NMT Model Size
5.7 Additional Analysis
We conduct additional analysis, including efficiency analysis (Appendix D.1), the case study on MGSM (Appendix D.2), and evaluation on multilingual commonsense reasoning (Appendix D.3).
@@ Appendix C Detailed Results
Appendix D Additional Analysis
@@ D.3 Multilingual Commonsense Reasoning
Af Akrikaans Id Indonesian Nl Dutch
Ar Arabic It Italian Pl Polish
Az Azerbaijani Ka Georgian Ps Pashto
Cs Czech Kk Kazakh Pt Portuguese
De German Km Khmer Ro Romanian
@@
Table 3: The included 42 untuned languages and corresponding language codes.
@@
Avg 38.71 17.68 36.28 34.59 30.82 9.21 21.09 26.66
Table 11: Detailed results on 42 untuned languages, composing MetaMath-7B with NLLB-200-1.3B. We evaluate cross-model mapping quality on FLORES-101, following the main experiments. "X" denotes all languages except for English. We bold the best scores for the LLM group.

### 16 Apr 2026 v4

@@ 2.1 Data-Level Multilingual Enhancement for LLMs
A line of work augments the multilingual capabilities of LLMs at the data level by constructing multilingual training corpora using pretrained multilingual or machine translation models Li et al. (2023); Zhang et al. (2023, 2024a, 2024b). Typical approaches translate English instruction into multiple languages Chen et al. (2024), pre-translate non-English inputs into English before task execution Qin et al. (2023); Chai et al. (2025), or leverage Mix-of-Experts (MoE) for language expansion Zhang et al. (2025b). Such approaches generally require continual multilingual training of LLMs, which may introduce translation noise and interfere with existing language capabilities. In practice, balancing performance across high- and low-resource languages remains challenging, as gains on low-resource languages often come at degradation on high-resource ones Gao et al. (2024). In contrast, XBridge achieves multilingual generalization through model composition without multilingual retraining of the LLM.
@@ 3.1 Architecture
Formally, given an input sequence $\mathbf{x}=(x_{1},\ldots,x_{n})$ in language $L_{x}$, we first encode it with the pretrained multilingual encoder $\mathtt{Enc}(\cdot)$, producing contextual representations $\mathbf{H}^{\mathbf{x}}\in\mathbb{R}^{n\times d_{e}}$. To bridge the representation gap between the multilingual encoder and LLM, we apply a lightweight mapping $\mathtt{Mapping_{enc}}(\cdot)$ that projects $\mathbf{H}^{\mathbf{x}}$ into the LLM representation space, yielding $\tilde{\mathbf{H}}^{\mathbf{x}}\in\mathbb{R}^{n\times d_{l}}$. The mapped encoder representations are then injected into the LLM together with a high-resource (English) instruction prompt, enabling the LLM to perform general knowledge processing conditioned on encoder semantics. Let $\mathbf{z}=(z_{1},\ldots,z_{m})$ denote the sequence of English tokens generated by the LLM. Rather than using the final-layer hidden states, we extract the penultimate-layer hidden states, denoted as $\mathbf{H}^{\mathbf{z^{\prime}}}\in\mathbb{R}^{m\times d_{l}}$, as Zhang et al. (2025a) show that the last layer is often tightly aligned with the output vocabulary space, while non-final layers retain richer semantic information.
@@ 5.2 Generalization to Untuned Languages
To examine whether the cross-model mappings learned by XBridge are language-agnostic rather than simply tied to specific training languages, we evaluate cross-model cross-lingual transfer on 41 untuned languages (listed in Table 3) in Figure 5, based on variants of Section 5.1.
@@ 5.6 Impact of NMT Model Size
5.7 Supplementary Analysis
We conduct supplementary analysis, including efficiency analysis (Appendix D.1), the case study on MGSM (Appendix D.2), and evaluation on multilingual commonsense reasoning (Appendix D.3).
@@ Appendix C Detailed Results
Appendix D Supplementary Analysis
@@ D.3 Multilingual Commonsense Reasoning
Af Akrikaans It Italian Pl Polish
Ar Arabic Ka Georgian Ps Pashto
Az Azerbaijani Kk Kazakh Pt Portuguese
Cs Czech Km Khmer Ro Romanian
@@
Id Indonesian Nl Dutch
Table 3: The included 41 untuned languages and corresponding language codes.
@@
Avg 38.71 17.29 36.28 34.57 30.78 8.98 21.10 26.64
Table 11: Detailed results on 41 untuned languages, composing MetaMath-7B with NLLB-200-1.3B. We evaluate cross-model mapping quality on FLORES-101, following the main experiments. "X" denotes all languages except for English. We bold the best scores for the LLM group.